OpenStudy (anonymous):

heyy help please? find the discriminant of the following equation x^2+2x+8=0

OpenStudy (anonymous):

x+2(2)(8) is this how it looks before simplifying?

OpenStudy (anonymous):

D=b^2-4ac=4-4*1*8=-28